You decide to travel by car for your holiday visits this year. You leave early in the morning to avoid congestion on the roads. This enables you to drive at a comfortable speed of 65.6 mph for 1.70 hours. However, after this time, you unexpectedly come to a stop for 23.4 min. Traffic starts moving again and you finish your travel at 57.4 mph for an additional 1.10 hours. How far did you travel on this trip? What was your average speed? There are 1609 meters in one mile.
Answer #1

Before the unexpected stop I drive at a speed 65.5 mph for 1.7 hours. So, during this time total distance covered by me is, d1 = 65.5 * 1.7 = 111.35 miles.

Then I have to stop for 23.4 mins = (23.4/60) hours = 0.39 hours.

After that delay I start again and end my journey with a speed of 57.4 mph for 1.1 hours. So during this interval I travel a distance of d2 = 57.4 * 1.1 = 63.14 miles.

Therefore during this wholr trip I travel a distance of d = d1 + d2 = 111.35 + 63.14 = 174.49 miles. Total time required during this trip is = 1.7 + 0.39 + 1.1 = 3.19 hours. Therefore the average speed is,

v = \frac{\text{total distance covered}}{\text{total time required to complete thie trip}} = \frac{174.49}{3.19} = 54.7 \hspace{2mm} \text{mph}